1.Let f(x)=x+1,g(x)=x^2-3. Solve the following equation for x [(fog)(x)+(gof)(x)=0]
2.Determine whether the following function is invertible or not, if it is , find its inverse,
g(x)=x-1/3x+2
1
Expert's answer
2020-11-03T18:17:12-0500

1. "fog(x) =f(g(x)) =(x^2-3)+1=x^2-2, \\
gof(x) =g(f(x)) =(x+1)^2-3=x^2+2x-2."

So

"fog(x) +gof(x) =2x^2+2x-4=0.
 \\
x^2+x-2=0, 
\\
x_1=-2, x_2=1."

Therefore, there are two roots of this equation.


2."\;g(x) =x-\dfrac{1}{3x}+2"

The invertible function should be strictly monotonic, so it cannot have the same value for different x. So it should have no more than 1 root. Let us determine the roots of g(x)

"0=x-\dfrac{1}{3x}+2, \\
0=3x^2+6x-1.
\\
D=b^2-4ac =36+12=48>0"


Therefore, the function has two roots and it can't be monotonic and invertible.